### 15 Mar 2026 v2

# Surrogate-Assisted Targeted Learning for Nested Bridge Functionals under Administrative Censoring

Abstract
We study semiparametric estimation of the average treatment effect when the target functional belongs to the nested bridge class: identification requires integrating an outcome regression over a conditional intermediate distribution rather than inverse-weighting incomplete observations. This structure arises whenever a primary outcome is administratively censored for a non-negligible fraction of units while an earlier surrogate remains broadly available, and it generates a distinctive obstruction for standard first-order estimators: a one-step debiased machine-learning construction leaves a second-order cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate and that has no doubly-robust complement in the efficient influence function.
We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that resolves this obstruction via a two-stage nested fluctuation step, and establish three structural results: (i) under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; (ii) the standard DML one-step remainder contains $R_{SY}$ as an irreducible second-order term (Proposition 1); and (iii) the two-stage targeting step absorbs $R_{SY}$ into the efficient score without estimating $f_{S}$, yielding an asymptotically linear, doubly robust estimator under a single product-rate condition (Theorem 2). For clustered data, valid inference requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2); a leave-one-cluster-out jackknife applied to the full estimator restores near-nominal finite-sample coverage where the first-order sandwich interval undercovers.
Stepped-wedge cluster-randomized trials, in which calendar-driven rollout creates near-boundary observation probabilities for late-crossing clusters, provide the sharpest instance of this functional class and serve as the running example. Simulations confirm near-zero bias and jackknife coverage of 0.956–0.968 across all cluster sizes evaluated, against a backdrop where inverse-weighted competitors break down entirely at high censoring rates.
1 Introduction
A class of causal functionals in modern longitudinal studies shares the following semiparametric structure. The target parameter — an average treatment effect — is identified not by inverse-weighting incomplete observations, but by integrating an outcome regression over the conditional distribution of an intermediate variable. We call this a nested bridge functional: its efficient score has a cross-product structure that creates a specific failure mode for standard first-order estimators, and resolving that failure mode requires a targeting step that goes beyond ordinary debiased machine learning.
The observational setting that generates this structure is one where a primary endpoint requires substantial follow-up to realize, so that only units enrolled early enough have a fully matured observation by the time of analysis. A shorter-term surrogate is observed broadly across the study population regardless of enrollment timing. Rather than inverse-weighting the incomplete observations by $g_{\Delta}^{-1}$ — which becomes unstable whenever $g_{\Delta}$ concentrates near zero — one identifies the average treatment effect by integrating the complete-case outcome regression $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ over the treatment-specific conditional distribution of the surrogate:
$\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl(\mathbb{E}_{S\mid A=1,W,t}\bigl[\bar{Q}_{Y}(S,1,W,t)\bigr]-\mathbb{E}_{S\mid A=0,W,t}\bigl[\bar{Q}_{Y}(S,0,W,t)\bigr]\Bigr).$ (1)
This representation requires only that $\bar{Q}_{Y}$ be estimable on the observed sub-cohort $\{\Delta=1\}$, not that $g_{\Delta}$ be bounded away from zero. The price is a new semiparametric complication: the functional (1) is nested in the sense that the outer integration is over the conditional surrogate law $f_{S}$, and the estimator must marginalize over $f_{S}$ without placing it in a denominator.
The core semiparametric obstruction.
The nested structure of (1) creates a specific failure mode for standard debiased machine-learning (DML) estimators. A natural one-step construction estimates $\bar{Q}_{\mathrm{int}}(a,W,t)\equiv\int\bar{Q}_{Y}(s,a,W,t)\,f_{S}(s\mid a,W,t)\,d\mu(s)$ by plugging in $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ separately, leaving a second-order cross-product remainder
$R_{SY}\;=\;\sum_{a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (2)
Unlike the standard doubly-robust remainder, $R_{SY}$ has no complement in the efficient influence function: no nuisance parameter whose consistent estimation forces $R_{SY}=0$. DML cross-fitting eliminates first-order empirical-process terms but leaves $R_{SY}$ unchanged, since $R_{SY}$ is a $P_{0}$-expectation of a second-order product. Controlling $R_{SY}=o_{P}(J^{-1/2})$ therefore requires either a rate condition on $\hat{f}_{S}$ or an estimator that avoids placing $\hat{f}_{S}$ in the expansion. The SA-TMLE takes the second path: its two-stage nested fluctuation enforces $N^{-1}\sum D_{S,a}(O_{ijt})=0$, constraining $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ to a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$ over observed surrogate values, absorbing $R_{SY}$ without ever estimating $f_{S}$.
Motivating instance.
Stepped-wedge cluster-randomized trials (SW-CRTs) provide a sharp instance of this structure. Calendar-determined treatment rollout means late-crossing clusters accumulate few primary-outcome observations before the administrative end date, creating a design-induced near-boundary regime $g_{\Delta}\to 0$ for those clusters. A short-term surrogate observed at crossover is available for all clusters, making the surrogate-bridge framework directly applicable. We use the SW-CRT as the running example throughout; the framework is not tied to this design.
Limitations of standard approaches.
IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation, producing severe variance inflation as $g_{\Delta}\to 0$ and potentially failing to admit regular asymptotic approximations near the boundary. Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort and require correct specification of the secular time trend and censoring mechanism (Kasza et al., 2019).
Contributions.
This paper makes three contributions to semiparametric estimation theory, instantiated in the SW-CRT setting.
First, we identify the nested bridge functional class under surrogate-mediated administrative censoring and establish a G-computation representation (Theorem 1) that avoids inverse observation weights entirely. The representation replaces the unstable $g_{\Delta}^{-1}$ with a support positivity condition on the complete-case outcome regression, extending the scope of doubly robust estimation to near-boundary censoring regimes.
Second, we characterize the efficient-score geometry of this functional class (Section 3). Three structural results follow: the censoring mechanism contributes no separate tangent-space component under surrogate-mediated MAR (Lemma 1); the nested cross-product remainder $R_{SY}$ is an irreducible second-order obstruction with no doubly-robust complement (Proposition 1); and valid inference under clustered dependence requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2).
Third, we construct the SA-TMLE and prove that its two-stage nested fluctuation absorbs $R_{SY}$ without estimating $f_{S}$, achieving $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition (Theorem 2). For finite-sample inference, a leave-one-cluster-out jackknife applied to the full estimator captures the fluctuation-step variance that the first-order sandwich misses, restoring near-nominal coverage across the cluster-count range evaluated.
Related statistical work.
Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes but does not address the nested bridge structure or the $R_{SY}$ obstruction. The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses the complementary problem of rare (not delayed) primary outcomes; their identification strategies do not extend to the near-boundary censoring regime. Luedtke et al. (2017) construct a semiparametric surrogate-index estimator for i.i.d. data structurally related to ours, but their asymptotic linearity requires a uniform positivity lower bound on $g_{\Delta}$ that the present approach does not need. The DML framework (Chernozhukov et al., 2018) provides product-rate guarantees via sample splitting, but cross-fitting does not eliminate $R_{SY}$; this paper isolates that remainder (Proposition 1) and resolves it via a second targeting step. [^1] See Kennedy (2022) for a review of debiased ML. Cluster-level TMLE (Balzer et al., 2016, 2021) establishes clusterwise influence aggregation for group-randomized trials without the nested bridge structure studied here.
Section 2 defines the observed data structure and establishes non-parametric identification via the surrogate-bridge representation. Section 3 develops the semiparametric theory, including the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and Proposition 1 characterizing $R_{SY}$. Section 4 constructs the SA-TMLE. Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen coverage bound. Section 6 reports Monte Carlo experiments. Section 7 provides the design-calibrated EPT illustration.
2 Data Structure and Identification
We now formalize the observed-data structure and identification strategy for the general nested surrogate-bridge problem described in Section 1. The framework applies whenever a delayed primary outcome is administratively censored for a subset of units at analysis time and an earlier surrogate is broadly observed. We instantiate it throughout in the stepped-wedge cluster-randomized setting, which makes the positivity boundary concrete and provides the simulation and illustration target of later sections; the identification theorem and semiparametric theory of Sections 2.23 are stated for the general structure.
Consider a SW-CRT with $J$ clusters, $T$ calendar time steps, and $n_{j}$ individuals in cluster $j$ ($N=\sum_{j}n_{j}$ total). Each cluster crosses over once at a design-determined time $\tau_{j}\in\{2,\ldots,T\}$, so $A_{jt}=I(t\geq\tau_{j})$; crossover times are assigned uniformly at random and $A_{j1}=0$ for all $j$. The observed data tuple is
$O_{ijt}=\bigl(W_{ijt},\;A_{jt},\;S_{ijt},\;\Delta_{ijt},\;\Delta_{ijt}\cdot Y_{ijt}\bigr)\sim P_{0},$ (3)
where $W_{ijt}\in\mathbb{R}^{p}$ are baseline covariates; $S_{ijt}$ is the short-term surrogate; $\Delta_{ijt}\in\{0,1\}$ indicates whether the delayed primary outcome was observed by the administrative end date; and $Y_{ijt}$ is that outcome. The product $\Delta_{ijt}\cdot Y_{ijt}$ makes explicit that $Y$ is structurally missing for late-crossing clusters.
Let $\mathbf{U}=(U_{W},\,U_{A},\,b_{j},\,U_{S},\,U_{\Delta},\,U_{Y})$ denote the exogenous noise variables. The cluster-level random effect $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}_{b})$ is i.i.d. across clusters. All other components are individual-level noise terms assumed mutually independent across individuals and clusters.
The following system defines the endogenous variables in their causal time-ordering:
$W_{ijt}$ $=f_{W}(U_{W,ij}),$ (baseline covariates) (4)
$A_{jt}$ $=I(t\geq\tau_{j}),\quad\tau_{j}=f_{A}(U_{A,j}),$ (treatment; deterministic given $\tau_{j}$) (5)
$S_{ijt}$ $=f_{S}(W_{ijt},\,A_{jt},\,b_{j},\,t,\,U_{S,ijt}),$ (short-term surrogate) (6)
$\Delta_{ijt}$ $=f_{\Delta}(S_{ijt},\,A_{jt},\,W_{ijt},\,t,\,U_{\Delta,ijt}),$ (observation indicator) (7)
$Y_{ijt}$ $=f_{Y}(W_{ijt},\,A_{jt},\,S_{ijt},\,b_{j},\,t,\,U_{Y,ijt}).$ (delayed primary outcome) (8)
Three structural features deserve emphasis. First, $A_{jt}$ is deterministic given $\tau_{j}$; the treatment mechanism is known by design and $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is plugged in directly rather than estimated. Second, the cluster random effect $b_{j}$ enters both $f_{S}$ and $f_{Y}$, inducing the ICC among all observations within cluster $j$ (Remark 1). Third — and most critically — $Y$ does not appear in $f_{\Delta}$ and $U_{\Delta}$ does not appear in $f_{Y}$. This is the structural encoding of Assumption 3: conditional on $S$, the censoring indicator $\Delta$ is independent of the outcome $Y$. In DAG terms, there is no directed edge $Y_{ijt}\to\Delta_{ijt}$ (Figure 1).
[…]
Figure 1: Directed Acyclic Graph for the SW-CRT Structural Causal Model. Red circles: endogenous variables. Blue circles: exogenous and design variables. Violet arrows: cluster random effect $b_{j}$, the source of ICC. The critical structural feature is the absence of a directed edge $Y_{ijt}\!\to\!\Delta_{ijt}$, encoding Assumption 3: once $S_{ijt}$ is observed, the censoring probability depends on $S$ but not directly on the unobserved $Y$. The secular time trend $t$ acts as a common cause of all endogenous variables and must be adjusted for in estimation.
Remark 1 (Intra-cluster correlation).
The shared $b_{j}$ induces ICC $\sigma^{2}_{b}/(\sigma^{2}_{b}+\sigma^{2}_{\varepsilon})$; inference must use the cluster-robust sandwich of Section 3 and Super Learner cross-validation at the cluster level (Supplement, Section S3).
2.1 Potential Outcomes and the Causal Target
For $a\in\{0,1\}$, the causal ATE is $\Psi(P_{0})=\mathbb{E}[Y_{ijt}(1)-Y_{ijt}(0)]$, where the expectation is over the marginal distribution of $(W_{ijt},b_{j},t)$.
2.2 Non-Parametric Identification
Assumption 1 (Consistency).
For $A_{jt}=a$: $S_{ijt}=S_{ijt}(a)$ and $Y_{ijt}=Y_{ijt}(a)$.
Assumption 2 (Sequential Randomization).
$(Y(0),Y(1),S(0),S(1))\perp A_{jt}\mid W_{ijt},t$. This holds by design since $\tau_{j}$ is randomized independently of outcomes.
Assumption 3 (Surrogate-Mediated MAR).
$P(\Delta_{ijt}=1\mid Y_{ijt},S_{ijt},A_{jt},W_{ijt},t)=P(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ijt},t)$ (Rubin, 1976). Assumption 3 is most defensible when censoring is calendar-driven rather than health-driven; strong surrogates satisfy causal proximity ($S$ on the path $A\to S\to Y$), temporal availability, and biological stationarity of $\mathbb{E}[Y\mid S,A,W,t]$.
Assumption 4 (Support Positivity).
$P(\Delta_{ijt}=1\mid S_{ijt}=s,A_{jt}=a,W_{ijt}=w,t)\geq c>0$ on the relevant support. Unlike IPCW, the surrogate-bridge target functional does not contain $g_{\Delta}^{-1}$; this is a support condition for identifying $\mathbb{E}[Y\mid S,A,W,t]$ from $\{\Delta=1\}$, not an inverse-weighting regularity condition.
Assumption 5 (Design-Based Treatment Support).
$0<\mathbb{E}_{t}[g_{A}(1\mid W,t)]<1$. Under uniform crossover randomization $g_{A}(1\mid W,t)=(t-1)/(T-1)$, known by construction, so $g_{A}$ is plugged in directly rather than estimated.
Theorem 1 (Identification via a Surrogate Bridge under Support Positivity).
Under Assumptions 15, the causal ATE is identified by the longitudinal G-computation formula:
$\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl($ $\mathbb{E}_{S\mid A=1,W,t}\bigl[\mathbb{E}[Y\mid S,A=1,W,t,\Delta=1]\bigr]$
$-\;\mathbb{E}_{S\mid A=0,W,t}\bigl[\mathbb{E}[Y\mid S,A=0,W,t,\Delta=1]\bigr]\Bigr).$ (9)
Proof.
By Assumptions 12, $\mathbb{E}[Y(a)\mid S(a),W,t]=\mathbb{E}[Y\mid S,A=a,W,t]$. By Assumption 3 and 4, conditioning on $\{\Delta=1\}$ is valid: $\mathbb{E}[Y\mid S,A=a,W,t]=\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]\equiv\bar{Q}_{Y}(S,a,W,t)$. Integrating over $P(S\mid a,W,t)$ and marginalizing over $(W,t)$ yields (1).∎∎
3 Semiparametric Theory
We now study the surrogate-bridge estimand as a semiparametric functional. This perspective is central for two reasons. First, it reveals the efficient-score geometry induced by surrogate-mediated censoring and clustered dependence: the censoring mechanism contributes no separate tangent-space component, and the cluster structure requires summation rather than averaging of influence contributions. Second, it clarifies why a standard one-step debiased estimator is not sufficient for the nested bridge problem: the second-order remainder contains a cross-product term involving the conditional surrogate law that is not removed by ordinary cross-fitting. The resulting analysis motivates the specific two-stage targeting construction developed in Section 4.
3.1 EIC Decomposition
The individual-level EIC components, whose derivation is provided in Supplement, Section S2, are as follows. For treatment $a\in\{0,1\}$:
$D_{Y,a}(O_{ijt})$ $=\frac{I(A=a)\,\Delta}{g_{A}(a\mid W,t)\,g_{\Delta}(1\mid S,A,W,t)}\bigl(Y-\bar{Q}_{Y}(S,\,a,\,W,t)\bigr),$ (10)
$D_{S,a}(O_{ijt})$ $=\frac{I(A=a)}{g_{A}(a\mid W,t)}\bigl(\bar{Q}_{Y}(S,\,a,\,W,t)\;-\;\bar{Q}_{\mathrm{int}}(a,\,W,t)\bigr),$ (11)
$D_{W,a}(O_{ijt})$ $=\bar{Q}_{\mathrm{int}}(a,W,t)-\hat{\Psi}_{a},$ (12)
where $\hat{\Psi}_{a}=N^{-1}\sum_{j,i}\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t_{ij})$.
In Equation (11), both nuisance functions are evaluated at the fixed intervention value $a$ rather than the observed random variable $A$. Under the indicator $I(A=a)$, the two are numerically equal on the support of the data, but the fixed-argument form is essential: it makes explicit that $\bar{Q}_{Y}(S,a,W,t)$ and $\bar{Q}_{\mathrm{int}}(a,W,t)$ are counterfactual predictions under $\mathrm{do}(A=a)$, consistent with the identification formula (1). Writing $A$ in the argument of $D_{S,a}$ would obscure this causal interpretation and introduce an inconsistency with the identification strategy.
The total individual EIC is:
$D^{*}(O_{ijt})\;=\;\bigl(D_{Y,1}+D_{S,1}+D_{W,1}\bigr)\;-\;\bigl(D_{Y,0}+D_{S,0}+D_{W,0}\bigr).$ (13)
The first structural result shows that surrogate-mediated censoring changes the identification strategy but not the efficient-score geometry in the usual missingness direction: once the surrogate is conditioned upon, the censoring mechanism contributes no separate tangent-space component.
Lemma 1 (Vanishing $\mathcal{T}_{\Delta}$ Component).
Under Assumption 3 (Surrogate-Mediated MAR), the pathwise derivative of $\Psi_{a}$ in the $\mathcal{T}_{\Delta}$ direction is identically zero: the EIC has no censoring-mechanism component.
Proof.
Consider a parametric submodel perturbing only $g_{\Delta}$ along score $h_{\Delta}\in\mathcal{T}_{\Delta}$. The identification formula (Equation (1)) reaches $g_{\Delta}$ only through the restriction of $\mathbb{E}[Y\mid S,A{=}a,W,t]$ to $\{\Delta=1\}$. By Assumption 3, $P(\Delta=1\mid Y,S,A,W,t)=P(\Delta=1\mid S,A,W,t)$, so perturbing $g_{\Delta}$ while holding $\bar{Q}_{Y}(S,a,W,t)$ and $P(S\mid a,W,t)$ fixed leaves $\bar{Q}_{\mathrm{int}}(a,W,t)$ unchanged. Consequently,
$\frac{d}{d\epsilon}\Psi_{a}(P_{\epsilon})\big|_{\epsilon=0}=\mathbb{E}_{P}\!\bigl[\bigl(\bar{Q}_{\mathrm{int}}(a,W,t)-\Psi_{a}\bigr)\,h_{\Delta}(O)\bigr]=0,$ (14)
where the final equality holds because $\bar{Q}_{\mathrm{int}}(a,W,t)-\Psi_{a}$ is $P$-mean-zero and orthogonal to any score supported only on variations in the censoring mechanism. Without Assumption 3, the censoring direction contributes a non-zero term involving the unidentified quantity $\mathbb{E}[Y\mid S,A,W,t,\Delta=0]$, rendering the EIC non-estimable from observed data. ∎∎
The second structural result concerns the correct level of influence aggregation. Because the estimand is defined through cluster-randomized data with arbitrary within-cluster dependence, the efficient contribution of a cluster is obtained by summation rather than averaging of individual terms; this is essential for valid semiparametric variance estimation under unequal cluster sizes.
Lemma 2 (Cluster-Level Summation Rule).
Let $O_{j}=\{O_{ijt}:i=1,\ldots,n_{j},\;t=1,\ldots,T\}$ denote the full data vector for cluster $j$, and let $\Psi$ be expressed as a functional of the cluster-level distribution $P_{j}$. The Riesz representer of the pathwise derivative in $L^{2}(P_{j})$ — equivalently, the cluster-level EIC — is the sum (not average) of individual EICs:
$\mathrm{EIC}_{j}=\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt}).$ (15)
Averaging by $n_{j}T$ would rescale the influence curve, rendering the sandwich variance estimator inconsistent under unbalanced cluster sizes and non-negligible ICC.
Proof.
Collecting Proposition 3.1 for $a=1$ minus $a=0$:
$D^{*}(O_{ijt})\;=\;\bigl(D_{Y,1}+D_{S,1}+D_{W,1}\bigr)-\bigl(D_{Y,0}+D_{S,0}+D_{W,0}\bigr).$ (16)
Because clusters are the i.i.d. units of randomization, the parameter $\Psi$ is a functional of the cluster-level distribution $P_{j}$ of $O_{j}=\{O_{ijt}:i=1,\ldots,n_{j},\,t=1,\ldots,T\}$:
$\Psi=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{n_{j}}\bigl[\bar{Q}_{\mathrm{int}}(1,W_{ij},t_{ij})-\bar{Q}_{\mathrm{int}}(0,W_{ij},t_{ij})\bigr].$ (17)
Applying the pathwise derivative to a submodel $P_{j,\epsilon}$ perturbing only cluster $j$ along score $h_{j}(O_{j})$, the chain rule over the additive structure gives
$\frac{d}{d\epsilon}\Psi(P_{j,\epsilon})\big|_{\epsilon=0}=\mathbb{E}_{P_{j}}\!\!\left[\left(\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt})\right)h_{j}(O_{j})\right].$ (18)
The Riesz representer in $L^{2}(P_{j})$ is therefore the sum (not average) of individual EICs:
$\mathrm{EIC}_{j}\;=\;\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt}).$ (19)
Summation absorbs both the intra-cluster correlation (ICC) across individuals at a given step and the serial correlation across steps induced by the secular trend; averaging by $n_{j}$ would rescale the influence curve and produce inconsistent variance estimates under unbalanced cluster sizes. ∎
3.2 The Nested Cross-Product Remainder and the Necessity of Two-Stage Targeting
The previous two lemmas establish the efficient-score structure of the surrogate-bridge functional — the core semiparametric object of this paper — but they do not yet determine the appropriate estimator. A natural question is whether a standard one-step debiased machine-learning construction suffices. The next result shows that, for nested bridge functionals, the answer is generally no. Even after cross-fitting removes first-order empirical-process terms, a second-order cross-product remainder $R_{SY}$ remains because the bridge functional depends jointly on the outcome regression and the conditional surrogate law $f_{S}$. This remainder is not a fluctuation term and therefore is not eliminated by ordinary sample splitting. This is the semiparametric obstruction that the two-stage targeting step of Section 4 is specifically designed to resolve.
Proposition 1 (Nested Cross-Product Term in the DML Remainder).
Let $f_{S}(\cdot\mid a,W,t)\equiv P_{0}(S\in\cdot\mid A=a,W,t)$ denote the conditional density of the surrogate, and let $\hat{f}_{S}$ be any estimator of $f_{S}$ independent of the SA-TMLE nuisance estimators. Define the plug-in integral estimator $\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}}(a,W,t)\equiv\int\hat{\bar{Q}}_{Y}(s,a,W,t)\,\hat{f}_{S}(s\mid a,W,t)\,d\mu(s)$, which constructs $\bar{Q}_{\mathrm{int}}$ as an integral of $\hat{\bar{Q}}_{Y}$ against $\hat{f}_{S}$.
For the DML one-step estimator $\hat{\Psi}^{\mathrm{DML}}=\hat{\Psi}^{\mathrm{DML}}_{\mathrm{plug\text{-}in}}+N^{-1}\sum_{j,i,t}D^{*}(O_{ijt};\hat{\bar{Q}}_{Y},\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}},\hat{g}_{A},\hat{g}_{\Delta})$ based on this construction, the von Mises expansion yields:
$\hat{\Psi}^{\mathrm{DML}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;R_{2}^{\mathrm{TMLE}}(\hat{P},P_{0})\;+\;R_{SY}(\hat{P},P_{0})\;+\;o_{P}(J^{-1/2}),$ (20)
where $R_{2}^{\mathrm{TMLE}}$ is the TMLE remainder (Supplement, Section S2.4) and the nested cross-product remainder is
$R_{SY}(\hat{P},P_{0})\;=\;\sum_{a\in\{0,1\}}(-1)^{1-a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (21)
For $R_{SY}=o_{P}(J^{-1/2})$ the DML estimator requires the additional product-rate condition
$\sum_{a\in\{0,1\}}\bigl\|\hat{\bar{Q}}_{Y}(\cdot,a,\cdot,\cdot)-\bar{Q}^{0}_{Y}(\cdot,a,\cdot,\cdot)\bigr\|_{P_{0}}\cdot\bigl\|\hat{f}_{S}(\cdot\mid a,\cdot,\cdot)-f^{0}_{S}(\cdot\mid a,\cdot,\cdot)\bigr\|_{P_{0}}\;=\;o_{P}(J^{-1/2}),$ (22)
which involves estimating the conditional density $f_{S}$ at rate $o_{P}(J^{-1/4})$ in $L^{2}(P_{0})$.
The SA-TMLE eliminates $R_{SY}$ without estimating $f_{S}$: the second fluctuation step (Section 4.2) enforces $N^{-1}\sum_{j,i,t}D_{S,a}(O_{ijt})=0$, which constrains $\hat{\bar{Q}}^{\,*}_{\mathrm{int}}$ to the propensity-weighted empirical conditional mean of $\hat{\bar{Q}}^{\,*}_{Y}$, thereby absorbing $R_{SY}$ into the efficient score without requiring knowledge of $f_{S}$. The proof is given in Supplement, Section S2.6.
Remark 2 (Why DML cross-fitting does not eliminate $R_{SY}$).
[…]
4 Surrogate-Assisted TMLE Construction
Proposition 1 identifies the obstacle to a one-step debiased estimator: the nested bridge functional generates a second-order remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate. The estimator developed here resolves that obstacle directly. Rather than constructing the bridge by separately estimating the surrogate density, the proposed two-stage targeted minimum loss procedure absorbs $R_{SY}$ into the efficient score, yielding an estimator that preserves doubly robust asymptotic linearity without direct estimation of $f_{S}$.
4.1 Stage 1: Initial Estimation via Super Learner
We seek to estimate the population functional $\Psi(P_{0})$ defined in (1). We first construct initial estimates of the four nuisance functions via a Super Learner ensemble (Supplement, Section S3):
$\bar{Q}_{Y}(S,A,W,t)$ $=\mathbb{E}[Y\mid S,A,W,t,\Delta=1],$ (23)
$\bar{Q}_{\mathrm{int}}(A,W,t)$ $=\mathbb{E}\bigl[\bar{Q}_{Y}(S,A,W,t)\mid A,W,t\bigr],$ (24)
$g_{A}(1\mid W,t)$ $=P(A=1\mid W,t),$ (25)
$g_{\Delta}(1\mid S,A,W,t)$ $=P(\Delta=1\mid S,A,W,t).$ (26)
$\bar{Q}_{\mathrm{int}}$ is fit on the observed $A$ and evaluated counterfactually at $A=1$ and $A=0$ to produce $\hat{\bar{Q}}_{\mathrm{int}}(1,W,t)$ and $\hat{\bar{Q}}_{\mathrm{int}}(0,W,t)$; the TMLE fluctuation step corrects the resulting plug-in bias. The treatment propensity $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is plugged in directly from the randomization schedule; only $g_{\Delta}$ requires data-adaptive estimation.
4.2 Stage 2: The Nested Fluctuation Step
Clever covariate for the surrogate integration model.
$H_{S}(O_{ijt})\;=\;\frac{1}{\hat{g}_{A}(A_{\mathrm{obs}}\mid W,t)},$ (27)
with the second fluctuation update fit on the full data:
$\operatorname{logit}\!\bigl(\bar{Q}^{\,*}_{\mathrm{int}}(\varepsilon_{S})\bigr)\;=\;\operatorname{logit}\!\bigl(\hat{\bar{Q}}_{\mathrm{int}}(A_{\mathrm{obs}},W,t)\bigr)\;+\;\varepsilon_{S}\,H_{S}(O_{ijt}).$ (28)
Convergence criterion.
The fluctuation step solves the efficient score equation. We declare convergence when the empirical mean of the cluster-level EIC (Section 3) falls below the tolerance:
$\left|\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\right|\;\leq\;\frac{1}{\sqrt{N}\,\log N}.$ (29)
In practice a single one-step fluctuation is sufficient when the Super Learner initial estimates are well-specified; we iterate only if (29) is not met after the first step.
Point estimator.
The final TMLE point estimate is the empirical mean of the doubly-targeted counterfactual predictions:
$\hat{\Psi}_{\mathrm{TMLE}}\;=\;\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{n_{j}}\Bigl[\bar{Q}^{\,*}_{\mathrm{int}}(1,W_{ijt},t_{ij})\;-\;\bar{Q}^{\,*}_{\mathrm{int}}(0,W_{ijt},t_{ij})\Bigr].$ (30)
5 Asymptotic Theory
5.1 Cluster-Level Aggregation and Inference
We aggregate the individual influence curves to the cluster level by summing over all individuals and all time steps within cluster $j$:
$\mathrm{EIC}_{j}\;=\;\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt}).$ (31)
The double summation over both $i$ and $t$ is essential. Because the cluster is the independent unit of the data-generating process, $\mathrm{EIC}_{j}$ must absorb all within-cluster dependence: both the ICC across individuals at a given time step and the serial correlation across time periods induced by the secular trend. The sandwich variance estimator below is then valid without any additional autocorrelation correction, provided $J$ is sufficient for the cluster-level CLT. The formal projection argument establishing that summation (not averaging) is the correct operation is provided in Supplement, Section S2.
The cluster-robust sandwich variance estimator is:
$\widehat{\mathrm{Var}}\!\left(\hat{\Psi}_{\mathrm{TMLE}}\right)\;=\;\frac{1}{J}\cdot\frac{1}{J-1}\sum_{j=1}^{J}\left(\mathrm{EIC}_{j}-\overline{\mathrm{EIC}}\right)^{2},$ (32)
and a Wald confidence interval at level $1-\alpha$ is
$\mathrm{CI}^{z}_{1-\alpha}\;=\;\hat{\Psi}_{\mathrm{TMLE}}\;\pm\;z_{\alpha/2}\,\sqrt{\widehat{\mathrm{Var}}\!\left(\hat{\Psi}_{\mathrm{TMLE}}\right)},$ (33)
where $z_{\alpha/2}$ is the upper $\alpha/2$ quantile of $\mathcal{N}(0,1)$. For the small-$J$ regime, Theorem S1 (Supplement, Section S1) establishes that replacing $z_{\alpha/2}$ by the $t_{J-1}$ quantile strictly reduces the undercoverage risk; we denote this interval
$\mathrm{CI}^{t}_{1-\alpha}(J)\;=\;\hat{\Psi}_{\mathrm{TMLE}}\;\pm\;t_{J-1,\,\alpha/2}\,\sqrt{\widehat{\mathrm{Var}}\!\left(\hat{\Psi}_{\mathrm{TMLE}}\right)},$ (34)
where $t_{J-1,\alpha/2}$ is the upper $\alpha/2$ quantile of Student’s $t$-distribution with J1J-1 degrees of freedom.
5.2 Asymptotic Normality
We now show that the proposed estimator solves the semiparametric estimation problem induced by the nested bridge functional: after the two-stage targeting step, the estimator is asymptotically linear with influence function equal to the cluster-level efficient score, and the remaining second-order term obeys the usual product-rate control. Asymptotic linearity implies that the cluster-robust sandwich variance estimator is consistent for the true sampling variance.
Theorem 2 (Asymptotic Linearity and Normality of the Surrogate-Assisted TMLE).
Suppose Assumptions 15 hold and the following regularity conditions are satisfied:
  1. (C1)
    (Cluster independence) The cluster-level data vectors $O_{j}=\{O_{ijt}:i=1,\ldots,n_{j},\;t=1,\ldots,T\}$ are independent across $j=1,\ldots,J$, with $n_{j}$ bounded above by a constant $\bar{n}<\infty$.
  2. (C2)
    (Bounded outcomes) $Y$, $S$, and all nuisance functions are uniformly bounded almost surely under $P_{0}$.
  3. (C3)
    (Product-rate condition) The nuisance estimators satisfy
    $\bigl\|\hat{\bar{Q}_{Y}}-\bar{Q}^{0}_{Y}\bigr\|_{P_{0}}\cdot\bigl\|\hat{g_{\Delta}}-g^{0}_{\Delta}\bigr\|_{P_{0}}\;+\;\bigl\|\hat{\bar{Q}_{\mathrm{int}}}-\bar{Q}_{\mathrm{int}}^{0}\bigr\|_{P_{0}}\cdot\bigl\|\hat{g_{A}}-g^{0}_{A}\bigr\|_{P_{0}}\;=\;o_{P}(J^{-1/2}),$ (35)
    where $\|\cdot\|_{P_{0}}$ denotes the $L^{2}(P_{0})$ norm. This condition is satisfied whenever each nuisance estimator converges at rate $o_{P}(J^{-1/4})$ in $L^{2}(P_{0})$.
Then the TMLE is asymptotically linear:
$\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;o_{P}(J^{-1/2}),$ (36)
and, as $J\to\infty$:
$\sqrt{J}\,\bigl(\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\bigr)\;\to_{d}\;\mathcal{N}\!\bigl(0,\;\sigma^{2}\bigr),\qquad\sigma^{2}\;=\;\mathrm{Var}(\mathrm{EIC}_{j}).$ (37)
The cluster-robust sandwich estimator (32) is consistent for $\sigma^{2}/J$, so the 95% confidence interval (33) achieves asymptotically exact coverage.
Proof.
The von Mises expansion of $\hat{\Psi}_{\mathrm{TMLE}}$ around $P_{0}$ yields
$\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;R_{2}(\hat{P},P_{0})\;+\;R_{SY}(\hat{P},P_{0})\;+\;o_{P}(J^{-1/2}),$ (38)
where $R_{2}$ is the standard bilinear nuisance-error remainder and $R_{SY}$ is the nested cross-product term of Proposition 1.
Elimination of $R_{SY}$. The second fluctuation step (Section 4.2) enforces the score equation $N^{-1}\sum_{j,i,t}D_{S,a}(O_{ijt})=0$ for each $a$. Expanding $D_{S,a}$ from (11) shows that this constraint sets the propensity-weighted empirical mean of $\bar{Q}^{\,*}_{Y}$ equal to $\bar{Q}^{\,*}_{\mathrm{int}}$ at the observed surrogate values, i.e.,
$\frac{1}{N}\sum_{j,i,t}\frac{I(A=a)}{\hat{g}_{A}(a\mid W,t)}\bigl[\bar{Q}^{\,*}_{Y}(S_{ijt},a,W_{ijt},t)-\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t)\bigr]\;=\;0.$ (39)
Because $\bar{Q}^{\,*}_{\mathrm{int}}$ is thereby defined as a propensity-weighted empirical average of $\bar{Q}^{\,*}_{Y}$ over the observed surrogate distribution—rather than as an explicit integral against an estimated $\hat{f}_{S}$—the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ that constitutes $R_{SY}$ never enters the expansion; $R_{SY}$ is absorbed into the efficient score without any requirement on the rate of $\hat{f}_{S}$. The formal calculation is in Supplement, Section S2.6.
Control of $R_{2}$. The residual remainder $R_{2}(\hat{P},P_{0})$ is bilinear in the nuisance errors $(\hat{Q}_{Y}-Q^{0}_{Y})$ and $(\hat{g}_{\Delta}-g^{0}_{\Delta})$, and in $(\hat{Q}_{\mathrm{int}}-Q^{0}_{\mathrm{int}})$ and $(\hat{g}_{A}-g^{0}_{A})$. Cauchy–Schwarz gives $|R_{2}|\leq\|\hat{Q}_{Y}-Q^{0}_{Y}\|_{P_{0}}\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}+\|\hat{Q}_{\mathrm{int}}-Q^{0}_{\mathrm{int}}\|_{P_{0}}\|\hat{g}_{A}-g^{0}_{A}\|_{P_{0}}$, which is $o_{P}(J^{-1/2})$ by condition (C3). When the CV-TMLE of Supplement, Section S3 is used, sample splitting eliminates the empirical process terms that would otherwise require a Donsker condition (Zheng and van der Laan, 2011).
Asymptotic normality. With $R_{2}+R_{SY}=o_{P}(J^{-1/2})$ established, the Lindeberg–Feller CLT applied to the i.i.d. cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$ gives (37). Consistency of the sandwich estimator (32) for $\sigma^{2}/J$ follows by the law of large numbers applied at the cluster level. The complete argument is in Supplement, Section S2.4.∎∎
Remark 3 (Double robustness and role of $R_{SY}$).
Condition (C3) is doubly robust (Bang and Robins, 2005): it holds if either the outcome models $(\bar{Q}_{Y},\bar{Q}_{\mathrm{int}})$ or the propensities $(g_{A},g_{\Delta})$ are consistently estimated. Since $g_{A}$ is known by design, (C3) reduces to $\|\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\|_{P_{0}}\cdot\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}=o_{P}(J^{-1/2})$. The key separation from DML is (39): the second fluctuation constrains $\bar{Q}^{\,*}_{\mathrm{int}}$ to a propensity-weighted mean of $\bar{Q}^{\,*}_{Y}$ over observed surrogates, so the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ constituting $R_{SY}$ never enters the expansion.
5.3 Non-asymptotic Coverage Bound
A Berry–Esseen bound (Theorem S1, Supplement) quantifies finite-sample undercoverage of the sandwich interval in terms of bounded cluster scores, variance concentration, and the nuisance remainder, giving a minimum cluster count $J^{*}\approx 27$ for 1% maximum undercoverage at the simulation DGP. Corollary S2 shows that the constant coverage shortfall in Block I is the finite-sample signature of estimation near the product-rate boundary: the fluctuation-step variance contributes an $O(J^{-1})$ term not captured by the sandwich, and this term does not shrink relative to the leading variance term under parametric nuisance rates.
The leave-one-cluster-out jackknife for the full SA-TMLE estimator resolves this gap in practice (Section 6.3), yielding near-nominal coverage at $J\geq 30$. Formal jackknife variance consistency for targeted estimators with data-adaptive nuisance follows from the arguments of Zheng and van der Laan (2011); a complete non-asymptotic treatment is deferred to future work.
6 Monte Carlo Simulation Study
Three simulation blocks examine the main implications of the theory: stability under increasing censoring (Blocks I, III), finite-sample double-robustness under nuisance misspecification (Block II), and coverage scaling with $J$ (Block I). Throughout, “double robustness” refers to asymptotic first-order identification; finite-sample performance depends on which nuisance component is misspecified and the magnitude of the residual second-order error.
6.1 Data-Generating Process
We simulate a SW-CRT with $J$ clusters, $T=7$ calendar time steps, and $n_{j}=40$ individuals per cluster per step ($N=J\times T\times n_{j}$). Crossover times $\tau_{j}$ are drawn uniformly from $\{2,\ldots,7\}$; administrative censoring is imposed structurally so that clusters crossing at steps T1T-1 and $T$ have near-zero $g_{\Delta}$ (outcome delay $t_{\mathrm{lag}}=2$, grace period $t_{\mathrm{grace}}=1$, overall censoring rate $\approx 28\%$). The cluster random effect satisfies $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.034)$ ($\mathrm{ICC}\approx 0.05$). Baseline covariates are $W_{ij,1}\sim\mathcal{N}(0,1)$, $W_{ij,2}\sim\mathrm{Bernoulli}(0.4)$, $W_{ij,3}\sim\mathrm{Uniform}(0,1)$. The secular trend is non-linear, $\lambda(t)=0.5\sin(\pi(t-1)/(T-1))+0.3((t-1)/(T-1))^{2}$, so that parametric linear-in-$t$ models are misspecified. The surrogate satisfies $S_{ijt}=0.8A_{jt}+0.4W_{ij,1}-0.3W_{ij,2}+0.5\lambda(t)+0.6b_{j}+\varepsilon^{S}_{ijt}$, $\varepsilon^{S}_{ijt}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.25)$; the coefficient $0.6$ on $b_{j}$ ensures $S$ is an informative bridge. The observation indicator and outcome follow
$\operatorname{logit}P(\Delta_{ijt}=1\mid S,A,W,t)$ $=3.0-2.5\cdot\mathbf{1}(t\geq T-1)-0.8\cdot\mathbf{1}(t=T)+0.4S_{ijt}+0.3A_{jt}-0.2W_{ij,3},$ (40)
$Y_{ijt}$ $=\Psi_{0}A_{jt}+0.5S_{ijt}+0.4W_{ij,1}-0.2W_{ij,2}+0.6W_{ij,3}+0.8\lambda(t)+b_{j}+\varepsilon^{Y}_{ijt},$ (41)
with $\varepsilon^{Y}_{ijt}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.64)$ and direct effect $\Psi_{0}=-0.28$. The true ATE is $\Psi^{*}=-0.28+0.5\times 0.8=0.12$, placed near the detection boundary for $J\approx 30$.
6.2 Simulation Scenarios
We organise the simulation study into four factorial scenario blocks. Each scenario is replicated 1,000 times.
Block I: Baseline performance across cluster counts.
The DGP is fixed at the specification above. We vary $J\in\{10,20,30,50,100\}$ at the baseline DGP ($\Psi^{*}=0.12$). This block assesses how finite-sample performance scales with the number of randomised clusters and whether the small-$J$ Lindeberg–Feller approximation is adequate for the cluster-level CLT.
Block II: Misspecified nuisance models.
$J=30$. Three scenarios probe Theorem 2’s double-robustness: (i) outcome models misspecified (non-linear trend and $S\times A$ interaction omitted), $g_{\Delta}$ correct; (ii) $g_{\Delta}$ misspecified (surrogate $S$ omitted from the censoring model), outcome models correct; (iii) both misspecified. Scenarios (i) and (iii) are expected to show bias; scenario (ii) tests the arm of double robustness where consistent outcome models compensate for $g_{\Delta}$ misspecification.
Block III: Increasing administrative censoring severity.
$J=30$, $\Psi^{*}=0.12$. We vary the administrative grace period $t_{\mathrm{grace}}\in\{0,1,2,3\}$, with $t_{\mathrm{grace}}=0$ representing strict administrative censoring (no grace period, so all clusters crossing over at the final two steps have $\Delta_{ijt}=0$ with probability approaching 1) and $t_{\mathrm{grace}}=3$ representing mild censoring. To ensure four distinct structural censoring regimes, we set $t_{\mathrm{lag}}=3$ for this block (versus the default $t_{\mathrm{lag}}=2$ in Blocks I, II, and IV). The overall censoring rates under the four values are approximately $\{43\%,28\%,16\%,8\%\}$. This block directly tests the theoretical prediction that IPCW variance explodes at high censoring rates while $\hat{\Psi}_{\mathrm{TMLE}}$ remains stable.
We compare three estimators: (i) GLMM (lme4, Gaussian family, linear-in-$t$ secular trend, cluster random intercept — deliberately misspecified in Blocks I and III–IV); (ii) IPCW, which upweights observed outcomes by the inverse estimated censoring propensity, with cluster-robust sandwich variance; and (iii) CV-TMLE (Supplement, Section S3, $V=10$ cluster-level folds), the recommended estimator. The treatment propensity $g_{A}$ is plugged in from the design for the TMLE variant.
6.3 Results
Performance is assessed over 1,000 replicates per scenario via bias, RMSE, 95% CI coverage, and power. Monte Carlo standard errors for coverage are $\approx 0.007$, so the nominal tolerance band is $[0.936,0.964]$.
Block I: Baseline performance across cluster counts.
Table S5 (Supplement) reports point-estimation metrics for all three estimators across $J\in\{10,20,30,50,100\}$; the inference comparison between sandwich and jackknife for CV-TMLE is in Table 2.
Point estimation. CV-TMLE achieves near-zero bias ($|\mathrm{Bias}|<0.004$) at all cluster counts, with RMSE improving steadily from 0.092 at $J=10$ to 0.030 at $J=100$. GLMM exhibits persistent bias of approximately +0.07+0.07 due to the misspecified secular time trend; its coverage collapses from 0.52 at $J=10$ to near zero at $J=100$ as the standard error shrinks around the biased point estimate. IPCW exhibits substantial positive bias (approximately +0.22+0.22) and markedly higher variance, consistent with the theoretical prediction that inverse weighting becomes unstable when $g_{\Delta}$ concentrates near zero for late-crossing clusters.
Sandwich variance and undercoverage. CV-TMLE with the sandwich interval achieves coverage of 0.87–0.89 across all $J$, well below the nominal 0.95. Table 1 diagnoses the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance contribution from the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which vary across leave-one-cluster-out fits. The empirical variance exceeds the mean sandwich by a factor of approximately 1.6 at all $J$; this ratio is stable because both the sandwich and the omitted term are $O(J^{-1})$ under the present nuisance specification, so neither diminishes faster than the other.
Jackknife variance and coverage restoration. The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), captures the fluctuation-step variance that the sandwich misses. Table 1 shows that the mean jackknife variance tracks the empirical variance closely at all $J$. The resulting coverage (Table 2, Jackknife columns) is near-nominal at the practically important cluster sizes: 0.956 at $J=30$ and 0.960 at $J=50$. At very small $J$ (10 and 20) the jackknife is mildly conservative (0.970–0.980), as the leave-one-out procedure loses a larger fraction of training data per fold and slightly overestimates variability; this is a well-understood property of the jackknife in small samples and represents a safe failure mode. By $J\geq 30$ the conservatism resolves and coverage is stable near 0.956–0.968 through $J=100$.
The power tradeoff is modest: at $J=30$, jackknife power is 0.57 versus sandwich power of 0.73, reflecting wider intervals. By $J=50$, jackknife power reaches 0.786, and by $J=100$ it is 0.970 — essentially the same as the sandwich at those sizes. The undercoverage was a variance-calibration issue, not a failure of the point estimator; the cluster jackknife resolves it without altering the bias or RMSE.
Table 1: Variance decomposition for CV-TMLE coverage (Block I, 500 replicates). $\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$. $\overline{\mathrm{Sand.}}$: mean sandwich variance estimate. $\overline{\mathrm{Jack.}}$: mean cluster-jackknife variance estimate. Ratio $\widehat{\mathrm{Var}}_{\mathrm{emp}}/\overline{\mathrm{Sand.}}$; values >1>1 indicate sandwich underestimation attributable to unaccounted fluctuation-step variance. The jackknife variance tracks the empirical variance closely, explaining the near-nominal jackknife coverage in Table 2.
$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) $\overline{\mathrm{Jack.}}$ ($\times 10^{-3}$) Ratio Sand. Cov. Jack. Cov.
10 8.44 4.80 9.39 1.76 0.878 0.980
20 4.83 2.83 5.37 1.71 0.869 0.970
30 2.93 1.84 3.23 1.59 0.892 0.956
50 1.80 1.11 2.01 1.62 0.866 0.960
100 0.87 0.54 0.97 1.61 0.878 0.968
Table 2: Block I CV-TMLE results: point estimation and inference comparison (500 replicates, $\Psi^{*}=0.12$, $T=7$, $n_{j}=40$). Bias and RMSE are identical across both variance estimators by construction. The jackknife is the leave-one-cluster-out procedure applied to the full two-stage estimator including both fluctuation steps. Full results for GLMM and IPCW comparators in Table S5 (Supplement). MC standard error for coverage $\approx 0.010$.
Sandwich Jackknife
$J$ Bias RMSE $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) Cov. Power Cov. Power
10 +0.002+0.002 0.090 8.15 0.878 0.410 0.980 0.204
20 0.001-0.001 0.068 4.66 0.869 0.554 0.970 0.332
30 +0.003+0.003 0.054 2.91 0.892 0.727 0.956 0.570
50 +0.003+0.003 0.043 1.80 0.866 0.897 0.960 0.786
100 0.001-0.001 0.029 0.85 0.878 0.997 0.968 0.970
Block II: Misspecified nuisance models.
Table 3 reports performance at $J=30$ under the three misspecification regimes of Section 6.2. Under scenario (ii) (outcome models correct, $g_{\Delta}$ misspecified), CV-TMLE achieves near-zero bias (<0.001<0.001) and coverage of 0.921, providing supportive finite-sample evidence for the asymptotic double-robustness result in the regime where the outcome regressions are correctly specified: the correctly specified outcome models drive the product-rate condition (35) toward $o_{P}(J^{-1/2})$, consistent with Remark 3.
Under scenario (i) (outcome models misspecified, $g_{\Delta}$ correct), CV-TMLE exhibits bias of 0.100 with coverage 0.426. This illustrates an important finite-sample limitation of the asymptotic double-robustness guarantee: when the misspecified outcome regression leaves a large structured error, correct specification of $g_{\Delta}$ alone may not be sufficient to deliver good finite-sample performance at $J=30$. In this setting, the targeting step appears to reduce but not eliminate the residual bias induced by the misspecified outcome model.
Under scenario (iii) (both sides misspecified), CV-TMLE is biased ($\mathrm{Bias}=0.092$) and under-covers (coverage 0.593), as expected when both nuisance components are misspecified, illustrating that the asymptotic double-robustness guarantee does not extend to simultaneous misspecification. The GLMM is biased identically in all three scenarios (bias =0.066=0.066, coverage =0.597=0.597) because it uses neither $g_{\Delta}$ nor the surrogate bridge. IPCW is biased in all scenarios due to the near-boundary observation probabilities for late-crossing clusters, with identical results in scenarios (ii) and (iii) where the same $g_{\Delta}$ misspecification applies.
Table 3: Block II simulation results: double robustness under nuisance misspecification ($J=30$, $\Psi^{*}=0.12$, 1,000 replicates). Scenario labels: (i) outcome models only misspecified; (ii) censoring propensity only misspecified; (iii) both misspecified.
Scenario Estimator Bias RMSE Coverage Power
(i): $\bar{Q}_{Y},\bar{Q}_{\mathrm{int}}$ misspec. GLMM +0.066+0.066 0.077 0.597 0.999
IPCW +0.219+0.219 0.224 0.787 0.939
CV-TMLE +0.100+0.100 0.113 0.426 0.990
(ii): $g_{\Delta}$ misspec. GLMM +0.066+0.066 0.077 0.597 0.999
IPCW +0.131+0.131 0.142 0.996 0.156
CV-TMLE +0.000+0.000 0.057 0.921 0.633
(iii): Both misspec. GLMM +0.066+0.066 0.077 0.597 0.999
IPCW +0.131+0.131 0.142 0.996 0.156
CV-TMLE +0.092+0.092 0.106 0.593 0.965
Block III: Bias amplification under increasing censoring.
Figure 2 displays estimator performance as censoring increases from 8% to 43%. Across the censoring range, CV-TMLE maintains substantially smaller bias than IPCW and GLMM, while its coverage declines from 0.92 to 0.77 as censoring becomes more severe (Panels A–B). This pattern is qualitatively consistent with the remainder-variance explanation from Block I: the point estimator remains comparatively stable, but the sandwich interval becomes increasingly incomplete as censoring amplifies the contribution of nuisance-estimation error. GLMM and IPCW coverage collapses to near zero. The near-100% rejection rates of GLMM and IPCW at high censoring (Panel C) are driven largely by bias rather than accurate signal recovery, as reflected in their simultaneous near-zero coverage. CV-TMLE power (0.65–0.73) reflects genuine signal detection.
Figure 2: Block III: bias (A), coverage (B), and power (C) under increasing censoring ($J=30$, $t_{\mathrm{lag}}=3$, 1,000 replicates). CV-TMLE bias stays near zero while GLMM/IPCW bias grows to 0.32. High GLMM/IPCW rejection rates at heavy censoring reflect bias, not signal (near-zero coverage).
ICC heterogeneity (Block IV).
Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the jackknife restores near-nominal coverage (0.946–0.968) across the full empirically observed ICC range in implementation-science SW-CRTs, extending the Block I finding to the ICC heterogeneity setting; full results are in Supplement, Section S4.
7 Design-Calibrated Illustration: Washington State EPT Trial
Setting.
The EPT trial (Golden et al., 2015) assessed whether a public health programme promoting free patient-delivered partner therapy could reduce chlamydia positivity among women aged 14–25 in Washington State. The four-wave stepped-wedge design randomized $J=23$ local health jurisdictions (LHJs) across $T=5$ time periods (one baseline, four post-baseline steps of six to eight months each): six LHJs crossed over at each of waves 1–3 and five at wave 4, with approximately 162 chlamydia tests per LHJ per period at sentinel clinical sites. The delayed primary outcome is 12-month chlamydia test positivity. This is administratively censored for late-crossing LHJs: wave-4 LHJs ($t_{j}=4$) have under 12 months of post-crossover follow-up, yielding an administrative censoring rate of 86% in that wave; wave-3 LHJs are partially censored (40%) from period 3 onward; overall 33.7% of records have missing 12-month outcomes. EPT uptake — whether the index patient accepted patient-delivered partner therapy within 30 days of STI diagnosis — is complete for all subjects regardless of wave and satisfies the three surrogate criteria of Assumption 3: causal proximity to the intervention, temporal availability before the 12-month outcome, and biological stationarity.
We generate a replication dataset from the SCM of Section 2, with parameters fixed to match the published summary statistics of Golden et al. (2015): control positivity 5.0%, ICC $\approx 0.05$ (cluster-level CV =0.0165=0.0165), N=18,630N=18{,}630 individual records, and a treatment effect of 10% relative reduction on the log-risk scale. The oracle ATE is therefore known by construction: $\Psi_{0}=-0.0039$ (risk difference), consistent with the original trial’s null finding. All calibration code is in the software repository.
Results.
All three estimators produce point estimates within 0.003 of the oracle ATE $\Psi_{0}=-0.0039$ and achieve coverage (Table 4; Figure 3). The substantive comparison is CI width: IPCW (width 0.068) is twice that of SA-TMLE (0.034), reflecting variance inflation from near-zero wave-4 censoring weights — precisely the near-boundary regime characterised by Proposition 1. The GLMM achieves the narrowest CI (0.026) at the cost of requiring correct model specification for validity.
Table 4: Oracle comparison for the design-calibrated EPT illustration. Data generated under the SCM of Section 2 with parameters matching Golden et al. (2015): $J=23$ clusters, $T=5$ periods, overall censoring 33.7%, wave-4 censoring 86.5%, oracle ATE =0.0039=-0.0039. The exercise assesses which estimator’s CI covers the known truth and at what cost in bias and CI width; it does not constitute an inferential claim about the EPT intervention.
Estimator ATE SE 95% CI Oracle covered?
Oracle ($\Psi_{0}$, known by construction) 0.0039-0.0039
GLMM (complete-case) 0.0012-0.0012 $0.0066$ (0.014, 0.012)(-0.014,\;0.012) Yes
IPCW 0.0059-0.0059 $0.0163$ (0.040, 0.028)(-0.040,\;0.028) Yes
SA-TMLE (proposed) 0.0008-0.0008 $0.0082$ (0.018, 0.016)(-0.018,\;0.016) Yes
Figure 3: Oracle comparison for the design-calibrated EPT illustration (calibrated to Golden et al. 2015). Horizontal bars are 95% cluster-robust confidence intervals; the vertical dashed line marks zero; the dotted line marks the known oracle $\Psi_{0}=-0.0039$. All three estimators cover the oracle, with point estimates within 0.003 of the truth. The key comparison is CI width: IPCW (width 0.068) is twice as wide as SA-TMLE (0.034), reflecting variance inflation from near-zero wave-4 censoring weights. GLMM achieves the narrowest CI (0.026) at the cost of model dependence. This figure illustrates oracle coverage under a calibrated design, not the causal effect of EPT.
8 Discussion
We developed a Surrogate-Assisted TMLE for the nested bridge functional $\Psi(P_{0})=\mathbb{E}_{W,t}[\mathbb{E}_{S|A=1}[\bar{Q}_{Y}(S,1,W,t)]-\mathbb{E}_{S|A=0}[\bar{Q}_{Y}(S,0,W,t)]]$ under administrative censoring. The key theoretical contribution is the identification and resolution of $R_{SY}$ (Proposition 1): a second-order remainder with no doubly-robust complement that makes a standard DML one-step insufficient for this functional class. The two-stage nested fluctuation absorbs $R_{SY}$ without estimating the conditional surrogate density $f_{S}$, recovering $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition. Simulations confirm near-zero bias and markedly lower variance than IPCW across all censoring regimes evaluated.
Finite-sample coverage.
The sandwich interval for CV-TMLE undercovers at 0.87–0.89 across all cluster counts. Table 1 identifies the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance of the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which contribute an additional $O(J^{-1})$ term that does not vanish faster than the leading sandwich term under the present nuisance specification. This is not a failure of the point estimator — bias is essentially zero at all $J$ — but a variance-calibration gap attributable to the two-stage targeting construction.
The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), resolves this gap by capturing the fluctuation-step variance the sandwich misses. Coverage is 0.956 at $J=30$ and 0.960–0.968 at $J\geq 50$ — near nominal across the practically important regime. At very small $J$ (10–20) the jackknife is mildly conservative (0.970–0.980); this is a well-understood property of leave-one-out procedures in small samples and represents a safe failure mode. We recommend the cluster jackknife as the default inference procedure for SA-TMLE, with the sandwich as a fast approximation when $J\geq 50$.
Assumption diagnostics and sensitivity.
Assumption 3 (Surrogate-Mediated MAR) is partially testable on the observed sub-cohort via the augmented regression $\mathbb{E}[Y\mid S,A,W,t,\hat{r}(A,W,t),\Delta=1]$, where $\hat{r}=\hat{P}(\Delta=1\mid A,W,t)/\hat{P}(\Delta=1\mid S,A,W,t)$; a zero coefficient on $\hat{r}$ is consistent with Assumption 3. When the assumption fails with violation $\gamma(S,A,W,t)\equiv\mathbb{E}[Y\mid S,A,W,t,\Delta=0]-\mathbb{E}[Y\mid S,A,W,t,\Delta=1]$, the leading-order bias is
$\mathrm{Bias}(\hat{\Psi}_{\mathrm{TMLE}})\;\approx\;\mathbb{E}\!\left[\tfrac{(1-g_{\Delta})}{g_{\Delta}}\bigl(I(A=1)\gamma(S,1,W,t)-I(A=0)\gamma(S,0,W,t)\bigr)\right].$ (42)
The tipping-point $\gamma^{*}$ solving $\hat{\Psi}_{\mathrm{TMLE}}=\widehat{\mathrm{Bias}}(\gamma^{*})$ extends Scharfstein et al. (1999) to the surrogate-mediated setting and is analogous to the $E$-value of VanderWeele and Ding (2017); we recommend reporting $\gamma^{*}$ routinely.
Scope and future work.
The surrogate-bridge framework applies wherever a delayed primary outcome is observed sparsely and an intermediate variable is broadly available; extension to other longitudinal designs beyond SW-CRTs is straightforward in principle. Two open directions are: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, requiring bracketing entropy bounds beyond those available for current Super Learner ensembles.
Supplementary Material
The Supplementary Material contains: Section S1, the Berry–Esseen coverage bound (Theorem S1 and Corollary S1); Section S2, the derivation of the efficient influence curve including the tangent space factorization, total EIC, cluster-robust variance estimator, second-order remainder and double robustness proof, and the DML cross-product remainder (Proposition 1 proof); Section S3, the Super Learner specification and cluster-level CV-TMLE implementation details.
All analyses were conducted in R (version 4.3 or later). The Surrogate-Assisted CV-TMLE estimator, including the cluster-level Super Learner wrapper, the nested fluctuation step, the cluster-robust sandwich variance estimator, and all four simulation scenario generators (Blocks I–III of Section 6, and the Block IV ICC simulations), are implemented in the open-source R package swcrtSurrTMLE, available at https://github.com/[author]/swcrtSurrTMLE. The package depends on SuperLearner (Polley and van der Laan, 2010) for ensemble estimation, lme4 for the GLMM comparator, and tmle for reference TMLE implementations. A fully reproducible script that replicates all tables and figures in this paper is included in the package vignette and archived on Zenodo at https://doi.org/10.5281/zenodo.[XXXXX] (DOI to be assigned upon acceptance).